OpenStudy (anonymous):

Solve the cube root of 2 minus 5 x equals 3

OpenStudy (anonymous):

\[\sqrt[3]{2-5x} = 3 \] first cube both sides: \[2-5x = 27\] so \[5x = 2-27 = -25\] and this should give \[x=-5\]

OpenStudy (anonymous):

Great explanation! thank you :)